Shift Summary 9 Aug., 1999 evening.

Nickolas Solomey (solomey@hep.uchicago.edu)
Mon, 9 Aug 1999 23:50:33 -0500

Shift Summary for day shift, 09 Aug., 99

On Shift: Herman and Nick

Hours of beam delivered at all intensities: 8.0/8.5
Hours of beam delivered at nominal intensity: 8.0/8.5
Hours of E832 data (special runs not included): 7.65/8.5
Hours of special runs (laser scans, muon runs...): 0 hrss

Accounting of lost time: 14 min changing tapes, several min. for kfrend,
and a series of no spills in the accelerator, 8 min with beam stop
closing itself, 8 min pipeline init problems.

========================================================================
Runs ended this shift: 14154 E832_beam
========================================================================
Tapes brought to FCC this shift: QKa24x (for a=C,S,K and x=3,4,5) QKB085
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Data taking
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14154 15:30 New shift crew takes over.
15:50 ktevl1a disk full alarm, we deleted LAS013007.dat
and LAS13009.dat both from May 18 before the run
started
247 Ktev ana errors, we ran kfrend.
315 Ktev ana errors, we ran kfrend.
14154 360 20:02 Tapes full, run ends.
14155 20:16 Run started.
1 20:17 Ktev ana errors, we ran kfrend.
15 Error (unusual never before seen) Reg_top/sem
is an order of magniture to high, went away
after one spill, that spill had a sem of 7.3e12.
32 Ktev ana errors, we ran kfrend.
51 No beam, SEM read normal but no beam to exp? odd!
52 Scaler errors out of range, or slow daq scaler
error we had a sdaq ktscaler not running error.
54 No none of those problems, our Beam-stop is
closed!
55 Called main control room, Max said our
beam stop closed maybe some switch-yard safety
signal closed it or an accident in MCR! He opened
it for us.
56 Good data taking again, the page of scaler alarms
went away, from DC current histogram we lost 5
spills due to a closed beam-stop.
69 Ktev ana erros, we ran kfrend.
70 No data taking, more ktevana errors, ran kfrend.
71 Beam stop snapped shut again!
72 MCR called, said there is a thunderstorm occuring
(we did not notice) and that there are multiple
glitches in the electrical grid.
74 Beam stop open, good data taking.
76 Acclerator abort at beginning of spill.
100 Pipeline error, but pipeline init froze during
process, had to clean up and then restart pipe.tcl
after which several kfrends.
106 good spill,data taking resumes.
23:53 Turned over to next shift a little early.


========================================================================
Action needing followed up on:
We labed 2 cases worth of split tapes, there are only 16 cases of
blank left in the hall (i.e. fugi cases are 20 per box so that is 320
tapes) so there must be more elsewhere? Someone who knows about the
necessary number of tapes might want to look into this.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey